OpenStudy (anonymous):

solve the homogeneous equation: (xy + y^2)dx - x^2 dy = 0 I know the answer is y = -x/(ln|x| + C) and y = 0

OpenStudy (jamesj):

Hence dy/dx = y/x + (y/x)^2 ------- (*) Set u = y/x. Then y = ux => y' = u + xu' Combining with (*) xu' + u = u + u^2 => xu' = u^2 => du/u^2 = dx/x => -1/u = ln|x| + C => u = -1/( ln|x| + C ) Ok, you see what to do now